OpenStudy (anonymous):

need help simplify this expression. ill give a medal. 10P6 10P4 these are my choices. A. 151,200 B. 420 C. 30 D. 5,040

OpenStudy (the_fizicx99):

10P^6/10P^4 ?

OpenStudy (anonymous):

Are these permutations or powers of 10?

OpenStudy (anonymous):

permutations

OpenStudy (anonymous):

please i need help

OpenStudy (the_fizicx99):

I'm afk, hold on.

OpenStudy (anonymous):

ok thank you

OpenStudy (the_fizicx99):

k back xD umm I've completely forgotten how to do permutations D:

OpenStudy (anonymous):

damn lol

OpenStudy (the_fizicx99):

xD @Darry&Gator is pro at math. Better than me. :3 (Don't let me down darry...)

OpenStudy (anonymous):

P^2 is the simplified version of this equation. :3

OpenStudy (the_fizicx99):

.__________. oh darry....

OpenStudy (anonymous):

huh.. im confused

OpenStudy (anonymous):

Ohmigosh, not one of the answer choice. Hold on a second. ~nervous giggles~

OpenStudy (anonymous):

thats not even close to any of my choices .. lol i was about to say

OpenStudy (anonymous):

Ohmigosh, my hand are sweaty. I can't do this sorry. Let me call some awesome people for you, though. @mathmale @ganeshie8 @agent0smith @mathstudent55 @ranga

OpenStudy (anonymous):

Sorry though. And this might help too. http://www.mathsisfun.com/combinatorics/combinations-permutations.html

OpenStudy (the_fizicx99):

LOL .____. you have failed me ;~; ..my excuse is i forgot, never liked them >.>

OpenStudy (anonymous):

lol l hate these to

OpenStudy (the_fizicx99):

Omg! you summoned my master (mathmale) :O

OpenStudy (anonymous):

OHMIGOSH REALLY?

OpenStudy (the_fizicx99):

Yesh :3

OpenStudy (anonymous):

He is the best. :3 Better than the other guy. [shh] ;o

OpenStudy (anonymous):

so can you guys still try to help me find my answer lol to this problem ..

OpenStudy (mathmale):

Howdy, guys, and thank you very much for the kind words! I didn't realize I was someone's "master." :) If I must be that, at least I'll try to be kind. Permutations are not among the topics I know and understand best. I hesitate to give out pat answers. However, my TI-83 Plus calculator has a command, nPr, which makes short work of the two different permutations you've expressed. 10 P 6 (6 items taken from 10) comes to 151200. 10 P 4 comes to 5040. I urge y ou to look up words such as "permutation" through Google searches. The material you find through Google varies in quality, of course. http://en.wikipedia.org/wiki/Permutation includes explanations of how to calculate permutations via formula instead of by calculator command.

OpenStudy (anonymous):

.. so my answer would be or is 151,200 or 5,040??

OpenStudy (anonymous):

Are these combinations? Are they permutations within combinations? What is the definition of 10P6? What is the formula?

OpenStudy (anonymous):

these are permutations.

OpenStudy (mathmale):

Regarding ".. so my answer would be or is 151,200 or 5,040?? " I interpret this problem as presenting two separate problems involving permutations. The answer to one is 151,200 and that to the other is 5,040. Again, please do an Internet / Google search for "permutation" and look for a formula (or formulas) for evaluation of permutations. http://en.wikipedia.org/wiki/Permutation

OpenStudy (mathmale):

If you, PP, happen to have a TI-83 or -84 calculator, I could guide you through the process of finding permutations on one of those calculators.

OpenStudy (anonymous):

Thanks for Wikipedia ref n P k = n!/(n-k)! and it is the number of different groups of k length when taken from n objects. 10P6 = 10!/(4!) = 10*9*8*7*6*5= 151,200.

OpenStudy (anonymous):

i dont have a calculater

OpenStudy (anonymous):

10P4 = 10!/(10-4)! = 10*9*8*7= 5040

OpenStudy (mathmale):

PP: Have you some other way to evaluate permutations and combinations, other than grinding out multiplication and division by hand?

OpenStudy (anonymous):

nope not really

OpenStudy (mathmale):

DWC: Please explain how you arrived at those results. I agree with your results; it's just that PP needs to know how to obtain them herself. Many thanks.

OpenStudy (anonymous):

so @douglaswinslowcooper im a guy mathmale.... :(

OpenStudy (anonymous):

my answer would be either A or D

OpenStudy (mathmale):

PP: If you have no calculator, then I'm afraid there's no alternative but to do the tedious multiplication and division by hand. DWC: DWC, to me, represents Douglas Winslow Cooper, and that sounds like a very masculine name, so fear not: I am not accusing you of being a girl!!

OpenStudy (mathmale):

As before, I believe you were given 2 separate problems in permutation to solve; A is the answer to the first one and D is the answer to the second one. But remember, you do need to learn how to evaluate these permutations on your own if you're going to be tested on this material.

OpenStudy (anonymous):

its not a seprate problem its all in the same problem..

OpenStudy (mathmale):

In cases like this, it's especially important to include the instructions that came with the problem. Are you positive that you've shared every bit of the problem statement here on OpenStudy?

OpenStudy (anonymous):

yea it says to simplify the expression

OpenStudy (anonymous):

i am really lost

OpenStudy (mathmale):

PP: I have a suggestion: photograph the original problem and upload the image to OpenStudy. Something's fishy here...I need to see the original problem.

OpenStudy (anonymous):

ok then one sec.

OpenStudy (anonymous):

i actaully tried to do that a day ago but it didnt work for some reason

OpenStudy (anonymous):

all it says is to simplify the expression

OpenStudy (mathmale):

PP: Are you able to act on my request (that you share an image of the original problem)?

OpenStudy (anonymous):

no it wont let me

OpenStudy (mathmale):

Have you downloaded a screen capture utility? These things let you capture any part of what's on your computer screen and share the resulting image with others. It's a tool worth having and using. perhaps one of the more technically savvy OpenStudy participants could help you find, download, install and use a screen capture utility.

OpenStudy (anonymous):

this computer is dumb i can draw it for you if you would like

OpenStudy (mathmale):

Certainly, if you have a cell phone camera and can photograph your drawing and share it. Take a look at the original problem and make absolutely certain that you've copying every bit of the instructions. You see, I interpret your original problem statement as including 2 separate problems, whereas you insist it's just one problem. That's due to ambiguity. So, i wanna see those instructions, every word and every letter of them.

OpenStudy (ranga):

I suspect the original problem is division. That is, 10P6 / 10P4 = ? If that is the case, the problem can be solved without a calculator or lengthy hand calculations. Just knowing the formula for permutations will do: Numerator: 10P6 = 10! / (10-6)! = 10! / 4! Denominator: 10P4 = 10! / (10 - 4)! = 10! / 6! Put them together: 10! / 4! / 10! / 6! Flip the denominator fraction and multiply: 10! / 4! * 6! / 10! = 6! / 4! = 6*5 = 30